# fMRI predictors based on language models of increasing complexity recover brain left lateralization

### v1

@@ 4 Discussion & Conclusion
By manipulating the size of artificial neural language models used to fit fMRI data of naturalistic language comprehension, we observed:
  1. 1.
    a linear relationship between the $r$-score (the cross-validated correlation between fMRI signal and the models’ predicted time-courses), and model size (the logarithm of number of parameters): see Fig. 2.
  2. 2.
    a breaking down of this relationship when focusing on the most complex models – with more than 3 billions parameters. Yet, in this range, indices assessing the performance of the model on NLP tasks are still predictive of brain score (see Fig. 5).
  3. 3.
    a left-right asymmetry emerging when encoding models are based on increasingly complex models. The relationship between asymmetry and model size (in log number of parameters) itself follows a linear relationship (see Fig. 6).
The first observation replicates the finding of Antonello et al., (2024) who first described a scaling law between model size and brain score, extending to brain data the scaling laws observed between model size and performance on various NLP tasks (Kaplan et al.,, 2020). Here, we generalize their result to a different dataset and using a larger variety of models and families that notably include a non-Transformer, Mamba.
The second point, concerning the relationship between neural network models performance in NLP tasks and brain scores, was also investigated by Schrimpf et al., (2021) and by Caucheteux and King, (2022). Both teams reported that brain scores generally improved as the performance in the next word prediction task increased, extending to the field of language processing results that were found in the visual domain by Schrimpf et al., (2020) or in the auditory domain by Kell et al., (2018). Caucheteux and King found that brain scores reached a plateau, and even dropped somewhat (see their Fig.2g), but we do not observe a plateau effect, which might be due to an overspecialization to the next word prediction task in their study. It should be noted that our large models are much bigger than those used in either studies, and have perplexities that are very low compared to them. It will be interesting for future work to consider even larger pretrained models to see if and when a plateau is reached.
The third point is the emergence of an asymmetry with increasing model size, an observation that has not been reported in the literature. A quick but vague explanation is that larger models better capture language representations in the left hemisphere. It begs the question of which representations are improved in larger models relative to smaller ones. For example, is it lexical, syntactic, semantic, pragmatic or world knowledge? The present work calls for detailed comparisons of the features discovered by large vs small language models, and how and where this translates into better fMRI prediction.
@@
Despite the asymmetries observed in other regions, it is remarkable that the largest models’ fits always improve over the smallest models even in the right hemisphere. It has been known for long that the right hemisphere has some language capacity (Bradshaw et al.,, 2017), which can manifest for example in split brain patients, or in patients with lesions in the left hemisphere. The right hemisphere is usually “hidden” because it is inhibited by the left hemisphere in healthy people (Vigneau et al.,, 2011). Speech prosody is usually considered to be mostly processed in the right temporal lobe (Wildgruber et al.,, 2006) and the right hemisphere seems to contribute to higher level aspects of language processing such as comprehending metaphors or jokes, detecting story inconsistencies (Jung-Beeman,, 2005; Bookheimer,, 2002).
An interesting follow-up for future work would be to see how the lateralization develops over time during the training process of a (very) large language model, either by training such a model from scratch or by having access to checkpoints during the training (see Antonello et al.,, 2024; Pasquiou et al.,, 2022, for first attempts). In humans, it seems that language processing in the brain starts bilaterally and then becomes lateralized over time (Szaflarski et al.,, 2006; Olulade et al.,, 2020; Ozernov-Palchik et al.,, 2024), although some studies have found very early lateralization (Witelson and Pallie,, 1973; Wood et al.,, 2004). Given our results on how better models recover the lateralization, we might expect the same to be true during the training process of a given model.
Limitations
  1. 1.
    Our analyses were performed at the group level, on data averaged over 48 right-handed participants. Performing similar analysis on individuals might be interesting to assess inter-individual variability. However, this would require a random sample of the population, not only right-handed participants, and, more crucially, an independent assessment of hemispheric dominance of these individuals, i.e. another measures of the asymmetry. In itself, we would not expect much novel insights from such an analysis.
  2. 2.
    We only analyzed data from English speakers. There is no reason to expect that our result is language-specific. From the neuro-linguistic point of view, left dominance is universal (Malik-Moraleda et al.,, 2022), and is even attested in sign language (Poizner et al.,, 1987). Nevertheless, it would be nice to generalize this inquiry to other languages. As Le Little Prince dataset comprises fMRI data from French & Chinese speakers, one could perform similar analyses on these language. Basically, this would require more pretrained models (and compute budget).
@@ Appendix A Supplementary Tables
Table A.1: List of all 28 large language models included in the study, grouped by family.
@@ Appendix B Supplementary Figures
Figure B.6: Analysis by Regions of Interest. Brain correlation as a function of the logarithm of the number of parameters, for all 28 models, for each ROI.

### v2

@@ 4 Discussion & Conclusion
By manipulating the size of artificial neural language models used to fit fMRI data of naturalistic language comprehension, we observed: (1) a linear relationship between the logarithm of the number of parameters of models and their ability to predict the fMRI time-courses (Fig. 2) and (2) a left-right asymmetry emerging when encoding models are based on increasingly complex models (Fig. 6).
The first observation replicates the finding of Antonello et al., (2024) who described a scaling law between model size and brain score, extending to brain data the scaling laws observed between model size and performance on various NLP tasks (Kaplan et al.,, 2020). Here, we generalize their result to a different dataset that includes three languages, and using a larger variety of models and families that notably include a non-Transformer, Mamba.
The scaling law breaks down when focusing on the largest models, above 3B parameters. Yet, in this range, indices assessing the performance of the model on NLP tasks are still predictive of brain score (Fig. 5). The relationship between neural network models performance in NLP tasks and brain scores, was also investigated by Schrimpf et al., (2021) and by Caucheteux and King, (2022). Both teams reported that brain scores generally improved as the performance in the next word prediction task increased, extending to the field of language processing results that were found in the visual domain by Schrimpf et al., (2020) or in the auditory domain by Kell et al., (2018). Caucheteux and King, (2022) found that brain scores reached a plateau, and even dropped somewhat (see their Fig.2g) which might be due to an overspecialization for the next word prediction task in their study (Caucheteux et al.,, 2023).
Our second observation is the emergence of an asymmetry with increasing model size, a finding that has not been reported in the literature  [^3] Note that the growing asymmetry with model size goes beyond showing a left-right difference in a single model, which could be explained by signal-to-noise ratio asymmetry (see, e.g., Caucheteux and King,, 2022, Fig. 2d). A quick but vague explanation is that larger models better capture language representations in the left hemisphere. It begs the question of which representations are improved in larger models relative to smaller ones. For example, is it lexical, syntactic, semantic, pragmatic or world knowledge? The present work calls for detailed comparisons of the features discovered by large vs small language models, and how and where this translates into better fMRI prediction.
@@
It is remarkable that in all regions of interest, even in the right hemisphere, the largest models’ fits always improve over the smallest models (see Fig.7). It has been known for long that the right hemisphere has some language capacity (Bradshaw et al.,, 2017), which can manifest in split brain patients or in patients with lesions in the left hemisphere (Vigneau et al.,, 2011). The right hemisphere has been associated with speech processing, especially prosody, usually considered to be processed in the right temporal lobe (Wildgruber et al.,, 2006), as well as with high-level aspects of language understanding involved in the comprehension of metaphors or jokes (Jung-Beeman,, 2005; Bookheimer,, 2002). More work will be needed to determine which aspects of LLMs allow them to fit these regions.
Our main analyses were performed on fMRI data averaged across participants because of computational power limitations. We report preliminary analyses on 5 participants showing that the emergence of asymmetry holds at the individual level (see Fig. B.13). Nevertheless there is an important variability between subjects in brain correlations, in the difference between left hemispheric vs. right hemispheric average brain correlations, and in the slope of the left-right difference. It would be interesting for future work to assess the reliability of these measures, and to compare them with independent assessments of individual language hemispheric dominance. Unfortunately, such measurements are not provided in the fMRI dataset that we use.
An interesting question concerns the evolution of lateralization during the training process of large language models, either by training such a model from scratch or by having access to checkpoints during the training (see Antonello et al.,, 2024; Pasquiou et al.,, 2022, for first attempts). In humans, it seems that language processing in the brain starts bilaterally and then becomes progressively lateralized over time (Szaflarski et al.,, 2006; Olulade et al.,, 2020; Ozernov-Palchik et al.,, 2024), although some studies have found very early lateralization (Witelson and Pallie,, 1973; Wood et al.,, 2004). Our preliminary exploration with the Pythia-6.9B model (Fig. B.10) reveals that the left-right asymmetry emerges during the training process. This leaves open the questions of which representations improved by learning are responsible for this asymmetry, and whether parallels with the development of language acquisition can be drawn.
5 Acknowledgments
We would like to sincerely thank all four anonymous reviewers, whose constructive feedback significantly contributed to the improvement of the paper. Many thanks also to Yair Lakretz, Emmanuel Chemla, and all the participants of the Linguae ML seminar at École Normale Supérieure.
@@ Appendix A Supplementary Tables
Table A.1: List of the 28 English large language models included in the study, grouped by family.
@@ Appendix B Supplementary Figures
a. Refer to caption
b. Refer to caption
c. Refer to caption
Figure B.1: Inter-subjects correlations (model free) in homologous cortical regions of the left and right hemispheres. (a) The 48 cortical areas from nilearn’s cort-maxprob-thr0-2mm version of the Harvard-Oxford atlas. (b) Relationship between right and left inter-subjects correlations (dots correspond to regions from the atlas) (c) Differences in correlations between left and right regions. The number of stars next to each bar indicates an estimate of the level of statistical significance, assessed with a two-sample t-test (*:$p<0.05$, **:$p<0.01$, ***:$p<0.001$).
@@
a. Refer to caption
b. Refer to caption
Figure B.7: Left-right difference in brain correlation as a function of the relative layer depth, split by family. Considering (a) the 25% most reliable voxels or (b) the whole brain volume.
a. Refer to caption b. Refer to caption
c. Refer to caption d. Refer to caption
Figure B.8: Influence of inter-subject correlations on brain scores and growth in asymmetry. (a) and (b): Same as Fig. 6, but brain correlations are normalized by inter-subject correlations. (c) Brain correlation (average over all models) as a function of the inter-subjects correlation in each of the 48 parcels of the Harvard-Oxford atlas depicted in Fig. B.1. (d) Slope in the relationship between the L-R difference in r-score and the logarithm of the number of parameters, as a function of inter-subjects correlation. Each square corresponds to a parcel in the Harvard-Oxford atlas.
Figure B.9: Analysis by Regions of Interest. Brain correlation as a function of the number of parameters (in log scale), for all 28 models, for each ROI, for each hemisphere. The horizontal dashed lines represent the mean inter-subjects correlation in each corresponding ROI (along with 95% confidence interval).
a. Refer to caption
b. Refer to caption
Figure B.10: Impact of model training. (a) Left-right difference in brain correlation as a function of the number of parameters (in log scale), for two model families (GPT-2 and Qwen1.5), before (gray dots) and after (blue dots) training. (b) Brain correlation (top) and left-right difference in brain correlation (bottom) as a function of the amount of training, for the Pythia-6.9b model. From the untrained model (0%) to the fully trained model (100% corresponds to one epoch through the 300B tokens training set). All results were computed on the 25% most reliable voxels.
a. Refer to caption
b. Refer to caption
c. Refer to caption d. Refer to caption e. Refer to caption
f. model name $n_{\text{parameters}}$ $n_{\text{layers}}$ $n_{\text{neurons}}$ ckiplab/gpt2-tiny-chinese 12M 4 312 ckiplab/gpt2-base-chinese 102M 12 312768 Qwen1.5-0.5B 464M 24 1024 Qwen1.5-1.8B 1.84B 24 2048 Qwen1.5-4B 3.95B 40 2560 Qwen1.5-7B 7.72B 32 4096 Qwen1.5-14B 14.17B 40 5120 Qwen2-0.5B 494M 24 896 Qwen2-1.5B 1.54B 28 1536 Qwen2-7B 7.62B 28 3584 h. Refer to caption
g. Refer to caption i. Refer to caption
Figure B.11: Results from Chinese participants and large language models trained on Chinese data. Analyses performed on all 35 Chinese participants from the multilingual fMRI dataset Le Petit Prince. The average subject has 26050 voxels. (a,b) Same as Fig. 1 (a,b). (c,d,e) Same as Fig. B.1 (a,b,c). (f) List of all large language models used in this Chinese study (pretrained with an autoregressive objective on Chinese language data), as available on the Hugging Face hub. (g,h) Same as Fig. 2 (a, b). (i) Same as Fig. 6 (b).
a. Refer to caption
b. Refer to caption
c. Refer to caption d. Refer to caption e. Refer to caption
f. model name $n_{\text{parameters}}$ $n_{\text{layers}}$ $n_{\text{neurons}}$ ClassCat/gpt2-base-french 124M 12 768 Qwen2-0.5B 494M 24 896 Qwen2-1.5B 1.54B 28 1536 Qwen2-7B 7.62B 28 3584 h. Refer to caption
g. Refer to caption i. Refer to caption
Figure B.12: Results from French participants and large language models trained on French data. Analyses performed on all 28 French participants from the multilingual fMRI dataset Le Petit Prince. The average subject has 27773 voxels. (a,b) Same as Fig. 1 (a,b). (c,d,e) Same as Fig. B.1 (a,b,c). (f) List of all large language models used in this French study (pretrained with an autoregressive objective on French language data), as available on the Hugging Face hub. (g,h) Same as Fig. 2 (a, b). (i) Same as Fig. 6 (b).
a. Refer to caption
b. Refer to caption
c. Refer to caption
d. Refer to caption
e. Refer to caption
Figure B.13: Individual analyses performed on five English participants. Each column corresponds to one participant, sub-EN057, sub-EN058, sub-EN059, sub-EN061 and sub-EN062 respectively. Nine large language models were used: all the models from the GPT-2 and Qwen1.5 families used in the main study (see Table A.1), ranging from 124M to 14.2B parameters. (a) For each individual, brain correlation on the whole brain volume as a function of the number of parameters (in log scale). (b) Same as (a), but for the left hemisphere (red) and for the right hemisphere (green). (c) Difference between left and right hemispheric brain correlations. (d) and (e) Same as (b) and (c) but considering the best 10% voxels of the left hemisphere vs. the best 10% voxels of the right hemisphere.